Density of a gas is found to be 5.46g/L at 27c and 2 bar pressure what will be tsk density at stp
 

Grade:11

1 Answers

Abhinav
6 years ago
We know that density =PM /RT Since d1×T1/P1 = d2×T2/P2 Therefore 5.76 × 300/2 = d2 × 273/1 So, d2 = 3 gm/dm^3